### v1

# Grothendieck’s proof of Hirzebruch-Riemann-Roch theorem

Abstract
The Riemann-Roch Theorem is one of the cornerstones of algebraic geometry, connecting algebraic data (sheaf cohomology) with geometric ones (intersection theory). This survey paper provides a self-contained introduction and a complete proof of the Hirzebruch-Riemann-Roch (HRR) Theorem for smooth projective varieties over an algebraically closed field.
Starting from the classical formulations for curves and surfaces, we introduce the two modern tools necessary for the generalization: the Grothendieck group $K_{0}(X)$ as the natural setting for the Euler characteristic, and the Chow ring $A_{\bullet}(X)$ as the setting for cycles and intersection theory. We then construct the fundamental bridge between these two worlds—the Chern character ($\mathrm{ch}$) and the Todd class ($\mathrm{td}$) —culminating in a full proof of the HRR formula:
$\chi(X,\mathcal{E})=\int_{X}\mathrm{ch}(\mathcal{E})\cdot\mathrm{td}(X)$
We conclude by showing how this general formula recovers the classical theorems for curves and surfaces.
1 Introduction
Given an algebraic variety $X$, we can adopt (at least) two perspectives in its study: the first one stems from the algebraic world, namely Morita’s theory, homological and commutative algebra which allow us to identify $X$ with its category of coherent $\mathcal{O}_{X}$-modules. On the other hand of course we have the geometry of $X$ which is usually phrased in terms of property of particular embeddings (e.g. in $\mathbb{P}^{n}$) or, and this is intrinsic in our variety, in terms of its subvarieties and their mutual relations, that is, in its intersection theory. The real interest in algebraic geometry comes from the interplay between these points of view and the Hirzebruch-Riemann-Roch theorem is one bright example of such a relation.
The quest for this kind of relations has been ultimately started by the work of Bernhard Riemann and Gustav Roch to address the problem of computing the dimension of the space of meromorphic functions on a compact Riemann surface with prescribed zeros and poles. On the one side one finds that, given a divisor DD [^1] At this point, a divisor may be thought of as a compact way of describing the zero (and polar) locus we’re interested in. Namely here $D=\sum_{\mbox{finite}}n_{i}P_{i}$ is a formal sum where $P_{i}$ are points and $n_{i}\in\mathbb{Z}$. In this case we simply define $\deg\left(D\right)=\sum n_{i}$., the relation is
$l\left(D\right)-l\left(K-D\right)=\deg\left(D\right)+1-g,$
where $l\left(D\right)$ is the dimension we were looking for, $K$ is a canonically defined divisor and $g$ is the genus (a topological invariant). Phrased in more modern terms, one can associate to every coherent $\mathcal{O}_{X}$-module $\mathcal{F}$ a sequence of finite dimensional vector spaces $H^{i}\left(X,\mathcal{F}\right)$ and, under this perspective, the integer $\ell\left(D\right)$ is just the dimension of $H^{0}\left(X,\mathcal{O}_{X}\left(D\right)\right)$. Working harder one can show that
[…]
this is a particular case of the Serre duality [^2] For a proof of Riemann-Roch theorem in these terms, you can look at [Har, Theorem IV.1.3, pag. 295] . At this point one is lead to guess how a generalization to higher dimensions looks like: the obvious generalization of the left hand side is given by the (finite) sum
[…]
this has the nice property of being additive on bounded exact sequences. The tough part is to understand the right hand side. The two-dimensional case, that is for smooth surfaces, was worked out essentially by Guido Castelnuovo and can be written in cohomological terms as [^3] For a proof, see [Har, Theorem V.1.6, pag. 362].
$\chi\left(S,\mathcal{O}_{S}\left(D\right)\right)=\frac{1}{2}D\cdot\left(D-K\right)+\chi\left(S,\mathcal{O}_{S}\right)$
where the $\cdot$ denotes the intersection product between divisors. This can be seen as an evidence that the right hand side has to do with the intersection theory of our variety, and this will be addressed in Section 3.
The main ideas are essentially two: in the first part we introduce the Grothendieck group of an exact category, this will be a universal group for “functions” which are additive on short exact sequences, therefore it is the right source both for $\chi\left(X,\bullet\right)$ and for the so-called Chern character $\mathrm{ch}\left(\bullet\right)$, on the other hand one can associate to $X$ a ring $A_{\bullet}\left(X\right)$ made up of classes of subvarieties modulo a suitable equivalence relation and whose product is given essentially by intersecting such varieties. One can wonder whether this can be further generalized, and the answer is yes, but at the cost of another strong idea, which is ultimately one of the cornerstones of Grothendieck’s approach to geometry: the focus of the theorem should be a morphism between two schemes rather than a scheme itself. This leads to the general Grothendieck-Riemann-Roch theorem and it is essentially the content of SGA6.
What we actually prove here is the so-called Hirzebruch-Riemann-Roch theorem, although this name usually refers to the complex case, where also transcendental methods are applied in the course of the proof.
Notation.
[…]
2 The $K$-theory
2.1 The Grothendieck group
Here we introduce the basics about the Grothendieck group of rings and schemes.
Definition 2.1.
A category with exact sequences is a full additive subcategory $\mathcal{P}$ of an abelian category $\mathcal{C}$ which is closed under extensions, that is, given an exact sequence
$0\to P^{\prime}\to P\to P^{\prime\prime}\to 0$
in $\mathcal{C}$ with $P^{\prime},P^{\prime\prime}$ in $\mathcal{P}$, then also $P$ is an object of $\mathcal{P}$. An exact sequence in $\mathcal{P}$ is an exact sequence in $\mathcal{C}$ involving only objects of $\mathcal{P}$. [^5] One should also add some smallness condition $\mathcal{P}$: in [Ros, Definition 3.1.1, pag.109] a category with exact sequences is supposed to have a small skeleton, that is $\mathcal{P}$ is equivalent to a small category $\mathcal{P}_{0}$ (i.e. a category whose objects form a set). We will ignore this kind of set theoretic issues and always suppose that our categories have small skeleton: in fact, as long as we’re concerned with category of modules over a scheme or over a (possibly non-commutative) ring, this condition is automatically satisfied (see [Ros, Examples 3.1.2 (2), (3), (7), pagg. 109-111], where he also deals with other categories of modules.
Example 2.2.
Our main examples of categories with exact sequences will be
  • An abelian category $\mathcal{C}$ is a category with exact sequences as long as it is equivalent to a small category.
  • For a ring $A$, let $\mathrm{Proj}_{A}^{fg}$ be the category of finitely generated projective $A$-modules and $\mathrm{Mod}_{A}^{fg}$ the category of finitely generated $A$-modules: they are categories with exact sequences, embedded in the category $\mathrm{Mod}_{A}$ of $A$-modules. When $A$ is noetherian, the category $\mathrm{Mod}_{A}^{fg}$ is itself abelian. Moreover $\mathrm{Proj}_{A}^{fg}$ has the remarkable property that every exact sequence is split.
  • For an algebraic variety $X$ (or, more generally, a scheme) the category $\mathrm{Coh}_{X}^{lf}$ of locally free coherent $\mathcal{O}_{X}$-modules, embedded in the category $\mathrm{QCoh}_{X}$ of quasi-coherent $\mathcal{O}_{X}$-modules or even in $\mathrm{Coh}_{X}$ of coherent $\mathcal{O}_{X}$-modules when $X$ is locally noetherian.
Definition 2.3.
Fix a category with exact sequences $\mathcal{P}$ and let $S_{\mathcal{P}}$ be the set of isomorphism classes of objects of $\mathcal{P}$. The Grothendieck group $K_{0}\left(\mathcal{P}\right)$ of $\mathcal{P}$ is defined as
$K_{0}\left(\mathcal{P}\right)=\frac{\mathbb{Z}\left[S_{\mathcal{P}}\right]}{R}$
where $\mathbb{Z}\left[S_{\mathcal{P}}\right]$ is the free abelian group on $S_{\mathcal{P}}$ and $R$ is the subgroup generated by the relation $\left[A\right]=\left[B\right]+\left[C\right]$ every time an exact sequence of the form
$0\to B\to A\to C\to 0$
exists in $\mathcal{P}$.
Example 2.4.
Let $A$ be a ring, we denote with $K_{0}\left(A\right)$ the group $K_{0}\left(\mathrm{\mathrm{Proj}}_{A}^{fg}\right)$ and by $G_{0}\left(A\right)$ the group $K_{0}\left(\mathrm{Mod}_{A}^{fg}\right)$. In the same vein, given a projective variety $X$, we denote with $K_{0}\left(X\right)$ the group $K_{0}\left(\mathrm{Coh}_{X}^{lf}\right)$ and $G_{0}\left(X\right)$ the group $K_{0}\left(\mathrm{Coh}_{X}\right)$ .
Remark 2.5.
Let $\mathcal{F}:\mathcal{C}\to\mathcal{D}$ be an exact functor between two categories with exact sequences, then one readily verifies that we have a well defined group homomorphism
$\mathcal{F}_{\ast}:K_{0}\left(\mathcal{C}\right)$ $\to$ $K_{0}\left(\mathcal{D}\right)$
$\left[X\right]$ $\mapsto$ $\left[\mathcal{F}\left(X\right)\right].$
Example 2.6.
Let $X$ be a projective variety over a field $k$, then we have a well defined homomorphism
$\chi\left(X,\bullet\right):G_{0}\left(X\right)$ $\to$ $\mathbb{Z}$
$\left[\mathcal{F}\right]$ $\mapsto$ $\sum_{i\geq 0}\left(-1\right)^{i}\dim_{k}H^{i}\left(X,\mathcal{F}\right)$
We see that the sum is actually finite in view, for example, of Grothendieck’s Vanishing [Har, Theorem III.2.7, page 208] and each term is finite-dimensional in view of [Har, Theorem III.5.2, page 228]. In order to see that this defines a group homomorphism, consider a short exact sequence in $\mathrm{Coh}_{X}$
$0\to\mathcal{F}\to\mathcal{G}\to\mathcal{H}\to 0$
and the associated long exact sequence of finite-dimensional $k$-vector spaces
$0\to H^{0}\left(X,\mathcal{F}\right)\to H^{0}\left(X,\mathcal{G}\right)\to H^{0}\left(X,\mathcal{H}\right)\to\dots\to H^{n}\left(X,\mathcal{H}\right)\to 0$
with $n=\dim X$. By additivity of the dimension on exact sequences we conclude that
$\chi\left(X,\mathcal{G}\right)=\chi\left(X,\mathcal{F}\right)+\chi\left(X,\mathcal{H}\right).$
Of course, exactly the same line shows that $\chi\left(X,\bullet\right)$ is defined as $K_{0}\left(X\right)\to\mathbb{Z}$.
Lemma 2.7.
Let $\mathcal{P}$ be a category with exact sequences seen as a full subcategory of an abelian category $\mathcal{C}$ and suppose that, for every epimorphism $\pi:X\to Y$ in $\mathcal{P}$, $\ker\left(\pi\right)$ is still an object of $\mathcal{P}$. Let
$0\to X_{n}\to\dots\to X_{1}\to X_{0}\to 0$
be an eact sequence in $\mathcal{P}$. Then the equality
$\sum_{i=0}^{n}\left(-1\right)^{i}\left[X_{i}\right]=0$
holds in $K_{0}\left(\mathcal{P}\right)$.
Proof.
Let us proceed by induction on $n$: the cases when $n\leq 2$ are trivial. Suppose the statement holds forn1n-1, then we can consider the two exact sequences
$0\to K\to X_{1}\to X_{0}\to 0$
$0\to X_{n}\to\dots\to X_{2}\to K\to 0$
and apply the inductive hypothesis. ∎
Lemma 2.8.
Let $\mathcal{C}$ be an abelian category and let $X_{\bullet}$ be a bounded complex in $\mathcal{C}$. Then we have the following equality in $K_{0}\left(\mathcal{C}\right)$
$\sum\left(-1\right)^{i}\left[X_{i}\right]=\sum\left(-1\right)^{i}\left[H_{i}\left(X_{\bullet}\right)\right].$
Proof.
Let $d_{i}:X_{i}\to X_{i-1}$ be the differentials, then we have exact sequences
$0\to\mathrm{Im}\left(d_{i+1}\right)\to\ker\left(d_{i}\right)\to H_{i}\left(X_{\bullet}\right)\to 0$
$0\to\ker\left(d_{i}\right)\to C_{i}\to\mathrm{Im}\left(d_{i}\right)\to 0.$
We compute
$\sum\left(-1\right)^{i}\left[H_{i}\left(X_{\bullet}\right)\right]$ $=$ $\sum\left(-1\right)^{i}\left[\ker\left(d_{i}\right)\right]-\sum\left(-1\right)^{i}\left[\mathrm{Im}\left(d_{i+1}\right)\right]$
$=$ $\sum\left(-1\right)^{i}\left[\ker\left(d_{i}\right)\right]+\sum\left(-1\right)^{i}\left[\mathrm{Im}\left(d_{i}\right)\right]$
$=$ $\sum\left(-1\right)^{i}\left[C_{i}\right].$
Proposition 2.9.
Let $A$ be a regular ring [^6] We say that a local ring $\left(A,\mathfrak{m},k\right)$ is a regular local ring if it is noetherian and $\dim A=\dim_{k}\frac{\mathfrak{m}}{\mathfrak{m}^{2}}$ We say that a (non necessarily local) ring $A$ is regular if it is noetherian and for every prime ideal $\mathfrak{p}\in\mathrm{Spec}\left(A\right)$, the local ring $A_{\mathfrak{p}}$ is regular. (resp. let $X$ be a smooth projective variety). Then the natural inclusions $K_{0}\left(A\right)\to G_{0}\left(A\right)$ and $K_{0}\left(X\right)\to G_{0}\left(X\right)$ are isomorphisms. [^7] Look at the Resolution Theorem [Ros, Theorem 3.1.13, pag. 121] (or the subsequent page for the $K_{1}$-analogue) for a generalization of this result. In general one has a pletora of dévissage theorems for reducing the study of a category to that of a well-behaved subcategory by means of suitable exact sequences (e.g. resolutions).
Proof.
Suppose $A$ is a regular ring and let $M$ be a finitely generated $A$-module, then, in view of Serre-Auslander-Buchsbaum’s Theorem (see for example [Rot, Theorem 8.62, page 491]) there exists an exact sequence
$0\to P_{n}\to\dots\to P_{0}\to M\to 0$
with $P_{i}$’s finitely generated and projective, hence the statement for $K_{0}\left(A\right)$ follows from Lemma 2.7. For smooth projective varieties the proof follows the same pattern, except that the existence of a finite resolution made up of locally free coherent sheaves is a bit more delicate and follows from Serre-Auslander-Buchsbaum’s Theorem combined with [Har, Exercise III.6.8, page 238]. ∎
From now on, when dealing with regular rings or varieties, we’ll always identify the corresponding $K_{0}$ and $G_{0}$.
Example 2.10.
  1. 1.
[…]
  2. 2.
    More generally, let $A$ be a regular local ring, then all finitely generated projective modules are free and hence they have a well-defined rank. Like in the case of fields, and in view of Proposition 2.9, we deduce that $K_{0}\left(A\right)\simeq\mathbb{Z}$ induced by the rank function.
Exercise 2.11.
Let $R$ be any commutative ring. Then the assignment
$\left(\left[P\right],\left[Q\right]\right)\mapsto\left[P\otimes_{R}Q\right]$
defines a commutative ring structure on $K_{0}\left(R\right)$. Make sure to spot where projectivity is used.
Proposition 2.12.
Let $R$ be a regular ring, then the assignment
$\left(\left[M\right],\left[N\right]\right)\mapsto\left[M\right]\cdot\left[N\right]=\sum_{i\geq 0}\left(-1\right)^{i}\left[\mathrm{Tor}_{i}^{R}\left(M,N\right)\right]$
makes $G_{0}\left(R\right)$ into a commutative ring. Moreover the inclusion $K_{0}\left(R\right)\to G_{0}\left(R\right)$ is a ring isomorphism (cfr. Proposition 2.9 and Exercise 2.11).
Proof.
In view of the explicit definition of the inverse map
$G_{0}\left(R\right)\to K_{0}\left(R\right),$
we see that requiring it to be a ring isomorphism amounts to the following: given two finitely generated $R$-modules $M,N$, choose two finite resolutions
$0\to P_{n}\to\dots\to P_{1}\to P_{0}\to M\to 0$
$0\to Q_{m}\to\dots\to Q_{1}\to Q_{0}\to N\to 0$
made up of finitely generated projective $R$-modules (they exist because the ring $R$ is supposed to be regular). Then we need to show that
$\sum_{i,j\geq 0}\left(-1\right)^{i+j}\left[P_{i}\otimes Q_{j}\right]=\sum_{k\geq 0}\left(-1\right)^{k}\left[\mathrm{Tor}_{k}^{R}\left(M,N\right)\right].$
Consider the bounded double complex $C_{\bullet.\bullet}$ with
$C_{i,j}=P_{i}\otimes Q_{j}.$
Then, in view of [Rot, Theorem 10.16, page 629], both spectral sequences associated to $C_{\bullet,\bullet}$ converge to the homology of $\mathrm{Tot}\left(C_{\bullet,\bullet}\right)_{\bullet}$ which is [^8] In fact, as th modules $P_{i}$’s and $Q_{j}$’s are all projective, all lines and columns are exact, therefore at the next page we find that the only non-zero row (or column) is given by this sequence of $\mathrm{Tor}$’s.
$H_{n}\left(\mathrm{Tot}\left(C_{\bullet,\bullet}\right)\right)=\mathrm{Tor}_{n}^{R}\left(M,N\right).$
In view of Lemma 2.8 we deduce our desidered equality. ∎
Remark 2.13.
I leave as an exercise to the willing reader to show directly that this definition really endows $G_{0}\left(R\right)$ with a ring structure. The only non trivial point to show is associativity, for which I know no proof (maybe using a spectral sequence for nested $\mathrm{Tor}$’s?).
Example 2.14.
Let $\mathcal{F}:\mathcal{C}\to\mathcal{D}$ be a left exact additive functor between two abelian categories and suppose that its right derived functors $\left\{R^{i}\mathcal{F}\right\}_{i\geq 0}$ exist and that, for every object $X$, we have $R^{i}\mathcal{F}\left(X\right)=0$ for $i\gg 0$ (depending on $X$). Then it is well defined a group homomorphism
$\mathcal{F}_{!}:K_{0}\left(\mathcal{C}\right)\to K_{0}\left(\mathcal{D}\right)$
given by
$\mathcal{F}_{!}\left[X\right]=\sum_{i\geq 0}\left(-1\right)^{i}\left[R^{i}\mathcal{F}\left(X\right)\right].$
Our main example will be the following: let $f:X\to Y$ be a morphism between projective varieties, then we define $f_{!}:K_{0}\left(X\right)\to K_{0}\left(Y\right)$ as $\left(f_{{}_{\ast}}\right)_{!}$. Note that $f$ has finite relative dimension, so that our assumptions are satisfied.
Exercise 2.15.
Let $f:A\to B$ be a (commutative) ring homomorphism. We have an induced homomorphism
$K_{0}\left(A\right)$ $\to K_{0}\left(B\right)$
$\left[P\right]$ $\mapsto\left[P\otimes_{A}B\right].$
Show that this works as long as we’re working with projective modules. What conditions are needed to extend this map to $G_{0}\left(A\right)\to G_{0}\left(B\right)$?
Proposition 2.16.
Let $f:X\to Y$ be a closed embedding between two projective varieties with $Y$ smooth. For a vector bundle $\mathcal{F}$ on $X$, let
$0\to\mathcal{F}_{r}\to\mathcal{F}_{r-1}\to\dots\to\mathcal{F}_{0}\to f_{\ast}\mathcal{F}\to 0$
be a locally free coherent resolution of $f_{\ast}\mathcal{F}$. Then the assignment
$f_{!}:K_{0}\left(X\right)$ $\to K_{0}\left(Y\right)$
is given by
$f_{!}\left[\mathcal{F}\right]=\sum_{i=0}^{r}\left(-1\right)^{i}\left[\mathcal{F}_{i}\right].$
Proof.
First, let us note that $R^{i}f_{\ast}=0$ for $i>0$ and then, as our varieties are projective, in view of [Har, Theorem III.8.8, b), pag. 252] the functor $f_{\ast}\left(\bullet\right)$ preserves coherence and we have a well-defined map
$G_{0}\left(X\right)$ $\to G_{0}\left(Y\right)$
$\left[M\right]$ $\mapsto\left[f_{\ast}M\right].$
Then $f_{!}$ is just the composition
$K_{0}\left(X\right)\hookrightarrow G_{0}\left(X\right)\to G_{0}\left(Y\right)\simeq K_{0}\left(Y\right),$
where the last isomorphism is the one in Proposition 2.9. ∎
2.2 An important example: the Grothendieck group of projective spaces
Lemma 2.17.
[…]
Proof.
In view of [Har, Theorem III.5.1, page 225] we have
[…]
as an1<0a-n-1<0. The same proof shows also the second statement. ∎
Theorem 2.18.
[…]
Proof.
[…]
$\phi\left(x^{t}\right)=\left[\mathcal{O}_{X}\left(t\right)\right].$
[…]
$0\to\mathcal{O}_{X}\to\bigoplus_{i=1}^{n+1}\mathcal{O}_{X}\left(1\right)\to\bigoplus_{i=1}^{{n+1\choose 2}}\mathcal{O}_{X}\left(2\right)\to\dots\to\mathcal{O}_{X}\left(n+1\right)\to 0.$
In view of Lemma 2.7, denoting with $\xi=\left[\mathcal{O}_{X}\left(1\right)\right]$, we have
$0$ $=$ $1-\left(n+1\right)\xi+{n+1\choose 2}\xi^{2}+\dots+\left(-1\right)^{n+1}\xi^{n+1}$
$=$ $\left(1-\xi\right)^{n+1}$
which shows in turn that $\phi$ factors thought
$\tilde{\phi}:\frac{\mathbb{Z}\left[x\right]}{\left(1-x\right)^{n+1}}\to K_{0}\left(X\right)$
and also that
$\left[\mathcal{O}_{X}\left(-1\right)\right]=\phi\left(\left(-1\right)^{n+1}\xi^{n}+\dots+{n+1\choose 2}\xi-\left(n+1\right)\right).$
In particular $\tilde{\phi}$ is surjective and $K_{0}\left(X\right)$ is generated by the classes
$\left[\mathcal{O}_{X}\right],\left[\mathcal{O}_{X}\left(1\right)\right],\dots,\left[\mathcal{O}_{X}\left(n\right)\right],$
that is, by $1,\xi,\dots,\xi^{n}$. Let
$\alpha=\sum_{i=0}^{n}a_{i}\xi^{i}=0$
with $\left(a_{0},\dots,a_{n}\right)\neq\left(0,\dots,0\right)$. Set $j=\max\left\{i\,\mid\,a_{i}\neq 0\right\}$, then
$0=\alpha\cdot\xi^{-j}=a_{0}\left[\mathcal{O}_{X}\left(-j\right)\right]+\dots+a_{j}\left[\mathcal{O}_{X}\right],$
but then Example 2.6 and Lemma 2.17 shows that
$0=\chi\left(X,\alpha\cdot\xi^{-j}\right)=a_{j}\cdot\chi\left(X,\mathcal{O}_{X}\right)=a_{j}$
which is a contradiction. ∎
3 Intersection theory
3.1 The Chow groups
Lemma 3.1.
Let $A$ be a local noetherian 1-dimensional domain with fraction field $K$. For any $a\in A\backslash\left\{0\right\}$, set
$\mathrm{ord}_{A}\left(a\right)=\ell_{A}\left(\frac{A}{aA}\right).$
Then $\mathrm{ord}_{A}$ gives a well-defined group homomorphism $K^{\ast}\to\mathbb{Z}$.
Proof.
Clearly we have
$\ell_{A}\left(\frac{A}{aA}\right)=\ell_{\frac{A}{aA}}\left(\frac{A}{aA}\right),$
hence we only need to see that $\frac{A}{aA}$ is an artinian ring. Since it is noetherian, we only need to see that it is 0-dimensional, but this is clear since it has only one prime ideal. Given $a,b\in A\backslash\left\{0\right\}$, we have an exact sequence
$0\to\frac{A}{aA}\to\frac{A}{abA}\to\frac{A}{bA}\to 0$
and hence
$\mathrm{ord}_{A}\left(ab\right)=\mathrm{ord}_{A}\left(a\right)+\mathrm{ord}_{A}\left(b\right)$
in view of the additivity of the length. It is just a formal fact now that the assignment
$K^{\ast}$ $\to$ $\mathbb{Z}$
$\frac{a}{b}$ $\mapsto$ $\mathrm{ord}_{A}\left(a\right)-\mathrm{ord}_{A}\left(b\right)$
is a group homomorphism. ∎
Definition 3.2.
Let $X$ be a projective variety, a $k$-cycle on $X$ is a finite (formal) sum
$\sum n_{i}\left[V_{i}\right]$
[…]
$\left(f\right)=\sum_{V\subseteq W}\mathrm{ord}_{V}\left(r\right)\cdot\left[V\right]$
where the sum is take over codimension one subvarieties $V$ of $W$ and $\mathrm{ord}_{V}$ is the order function associated to the 1-dimensional local ring $\mathcal{O}_{W,V}$ [^10] You may want to glance at [Har, Exercise I.3.13, page 22] (see Lemma 3.1). We say that a $k$-cycle $\alpha$ is rationally equivalent to 0, written $\alpha\sim 0$, if
$\alpha=\sum\left(f_{i}\right)$
[…]
$A_{k}\left(X\right)=\frac{Z_{k}\left(X\right)}{\mathrm{Rat}_{k}\left(X\right)}$
and
$A_{\bullet}\left(X\right)=\bigoplus_{k\geq 0}A_{k}\left(X\right).$
Example 3.3.
If $d=\dim X$, then clearly we have $A_{i}\left(X\right)=0$ for $i>d$ and, as $X$ is irreducible, we have
$A_{d}\left(X\right)=\mathbb{Z}\cdot\left[X\right]\simeq\mathbb{Z}.$
Remark 3.4.
[…]
$f_{\ast}\left[V\right]=\begin{cases}\deg\left(V\to f\left(V\right)\right)\cdot\left[f\left(V\right)\right]&\mbox{if $\dim V=\dim f\left(V\right)$}\\ 0&\mbox{otherwise}\end{cases},$
then $f_{\ast}$ gives a graded group homomorphism
$f_{\ast}:A_{\bullet}\left(X\right)\to A_{\bullet}\left(Y\right),$
i.e. it respects rational equivalence (see [Ful, Theorem 1.4, page 11]).
Lemma 3.5.
Let $X\overset{g}{\longrightarrow}Y\overset{f}{\longrightarrow}Z$ be morphisms between projective smooth varieties. Then
$\left(f\circ g\right)_{\ast}=f_{\ast}\circ g_{\ast}$
as maps $A_{\bullet}\left(X\right)\to A_{\bullet}\left(Z\right)$.
Proof.
Working case by case, one is ultimately lead to showing the following:
Claim.
Let $X$ be irreducible and let $f,g$ be finite morphisms, then $\deg\left(f\circ g\right)=\deg\left(f\right)\circ\deg\left(g\right)$.
which is equivalent to the multiplicativity of the degree in finite field extensions. ∎
Remark 3.6.
[…]
$f_{\ast}:A_{0}\left(X\right)$ $\to$ $A_{0}\left(\left\{\ast\right\}\right)=\mathbb{Z}$
$\sum n_{i}\left[P_{i}\right]$ $\mapsto$ $\sum n_{i}$
The map $\pi_{\ast}$ is usually denoted with
$\int_{X}:A_{\bullet}\left(X\right)\to\mathbb{Z}.$
3.1.1 Intersection with divisors
In this section, in view of what follows, we have to deal with possibly singular varieties, referring to [Ful] for most of the proofs.
Definition 3.7.
Let $X$ be a projective variety and $j:Y\to X$ be the embedding map of a $k$-dimensional irreducible subvariety. For a Cartier divisor $D\subseteq X$, set
$D\cdot Y:=j_{\ast}\left[j^{\ast}\mathcal{O}_{X}\left(D\right)\right]\in A_{k-1}\left(X\right)$
where $\left[j^{\ast}\mathcal{O}_{X}\left(D\right)\right]$ is the Weil divisor on $Y$ associated to the line bundle $j^{\ast}\mathcal{O}_{X}\left(D\right)$.
Remark 3.8.
Note that $D$ appears in the form of $\mathcal{O}_{X}\left(D\right)$, therefore we get a pairing
$\mathrm{Pic}\left(X\right)\otimes_{\mathbb{Z}}Z_{k}\left(X\right)\to A_{k-1}\left(X\right).$
One pleasant feature of the pairing just introduced comes from the following Proposition
Proposition 3.9.
Let $X$ be a $d$-dimensional projective variety, then
  1. 1.
    given two Cartier divisors $D,E\subseteq X$, denoting with $\left[D\right],\left[E\right]$ the corresponding Weil divisors, we have
    $D\cdot\left[E\right]=E\cdot\left[D\right]$
    in $A_{d-2}\left(X\right)$;
  2. 2.
    the pairing of Remark 3.8 induces a map
    $\mathrm{Pic}\left(X\right)\otimes_{\mathbb{Z}}A_{k}\left(X\right)\to A_{k-1}\left(X\right).$
Proof.
The proof of point $1)$ is quite involved, so we refer to [Ful, Theorem 2.4, pag. 35] for a detailed discussion. Once one believes that $1)$ is true, then $2)$ comes easily: given a $\left(k+1\right)$-dimensional subvariety $V\subseteq X$ and a rational function $f$ on $V$, then by part $1)$
$D\cdot\left[\left(f\right)\right]=\left(f\right)\cdot\left[D\right]=0$
since $\left(f\right)=0$ in $\mathrm{Pic}\left(V\right)$. ∎
3.2 Projective bundles
Definition 3.10.
Let $X$ be a projective variety and let $\mathcal{E}$ be a locally free coherent $\mathcal{O}_{X}$-module of rank $r+1$. We define $\mathbb{P}_{X}\left(\mathcal{E}\right)$ as the $\mathbb{P}^{r}$-fibration over $X$ with fibres defined as $\mathbb{P}_{X}\left(\mathcal{E}\right)_{x}=\mathbb{P}\left(\mathcal{E}_{x}\right)$ for every $x\in X$. We call $\mathbb{P}_{X}\left(\mathcal{E}\right)$ the projective bundle associated to $\mathcal{E}$. We call the natural map $\mathbb{P}_{X}\left(\mathcal{E}\right)\to X$ its structure map.
Remark 3.11.
There are many other ways to introduce $\mathbb{P}_{X}\left(\mathcal{E}\right)$, here we review two of them:
(Scheme-theoretic)
Since $\mathcal{E}$ is coherent, we have a quasi-coherent sheaf of $\mathcal{O}_{X}$-algebras $\mathrm{Sym}\left(\mathcal{E}\right)$ which is locally isomorphic to $\mathcal{O}_{X}\left[T_{0},\dots,T_{r}\right]$, therefore we can define $\mathbb{P}_{X}\left(\mathcal{E}\right)$ as the relative projective space
$\pi:\mathbb{P}_{X}\left(\mathcal{E}\right)=\underline{\mathrm{Proj}}_{X}\left(\mathrm{Sym}\left(\mathcal{E}\right)\right)\to X.$
In view of the previous remark, every point of $X$ has an affine open neighborhood
$U=\mathrm{Spec}\left(A\right)\subseteq X$
such that
$\mathbb{P}_{X}\left(\mathcal{E}\right)\times_{X}U\simeq\mathbb{P}_{A}^{r},$
whence the name. Moreover the various $\mathcal{O}_{\mathbb{P}_{A}^{r}}\left(1\right)$ glue together to give a canonical invertible sheaf $\mathcal{O}_{\mathbb{P}_{X}\left(\mathcal{E}\right)}\left(1\right)$ on $\mathbb{P}_{X}\left(\mathcal{E}\right)$. See [Har, Section II.7, pag. 160-162] for a more precise account and a universal property. Let us only point out that, in view of [Har, Proposition II.7.10, pag. 161], $\pi$ is a projective morphism whenever $X$ admits an ample line bundle, therefore, if $X$ is a projective variety, also $\mathbb{P}_{X}\left(\mathcal{E}\right)$ is.
(Bundle-theoretical)
We can see $\mathcal{E}$ as a vector bundle on $p:E\to X$, hence we have a covering $\left\{U_{i}\right\}_{i}$ of $X$ and isomorphisms $\psi_{i}:\pi^{-1}\left(U_{i}\right)\to U_{i}\times\mathbb{A}^{r+1}$ such that, for every $i\neq j$, the maps
$\psi_{i}\circ\psi_{j}^{-1}:\left(U_{i}\cap U_{j}\right)\times\mathbb{A}^{r+1}\to\left(U_{i}\cap U_{j}\right)\times\mathbb{A}^{r+1}$
are given by
$\left(P,x\right)\mapsto\left(P,M_{ij}\cdot x\right)$
for $M_{ij}\in\mathrm{GL}_{r+1}\left(\mathcal{O}_{X}\left(U_{i}\cap U_{j}\right)\right)$. To define $\mathbb{P}_{X}\left(\mathcal{E}\right)$, we glue the patches
$U_{i}\times\mathbb{P}^{r}\overset{p_{1}}{\longrightarrow}U_{i}$
along
$\left(U_{i}\cap U_{j}\right)\times\mathbb{P}^{r}$ $\to\left(U_{i}\cap U_{j}\right)\times\mathbb{P}^{r}$
$\left(P,x\right)$ $\mapsto\left(P,\overline{M_{ij}}\cdot x\right)$
where $\overline{M_{ij}}\in\mathrm{PGL}_{r+1}\left(\mathcal{O}_{X}\left(U_{i}\cap U_{j}\right)\right)$ denotes the class of $M_{ij}$.
Remark 3.12.
On $\mathbb{P}_{X}\left(\mathcal{E}\right)$ there is another “distinguished” sheaf, namely the pull-back $\pi^{\ast}\mathcal{E}$: it is locally free of rank $r+1$ (since $\mathcal{E}$ is) and a choice of a local basis amounts to the choice of local coordinates for $\mathbb{P}_{X}\left(\mathcal{E}\right)$, hence we have a natural map $\pi^{\ast}\mathcal{E}\to\mathcal{O}_{\mathbb{P}_{X}\left(\mathcal{E}\right)}\left(1\right)$ which is surjective as $\mathcal{O}_{\mathbb{P}_{X}\left(\mathcal{E}\right)}\left(1\right)$ is generated by the coordinates.
Definition 3.13.
Let $X$ be a projective variety, $\mathcal{E}$ a locally free coherent $\mathcal{O}_{X}$-module of rank $r+1$ and $\pi:\mathbb{P}_{X}\left(\mathcal{E}\right)\to X$ the associated projective bundle. Denote with $D_{\mathcal{E}}\subseteq\mathbb{P}_{X}\left(\mathcal{E}\right)$ the Cartier divisor corresponding to $\mathcal{O}_{\mathbb{P}_{X}\left(\mathcal{E}\right)}\left(1\right)$. For every $i\geq-r$, set
$s_{i}\left(\mathcal{E}\right):A_{k}\left(X\right)$ $\to A_{k-i}\left(X\right)$
$\alpha$ $\mapsto\pi_{\ast}\left(\left(r+i\right)D_{\mathcal{E}}\cdot\pi^{\ast}\alpha\right),$
called the $i$-th Segre class.
Proposition 3.14.
Let $X$ be a projective variety:
  1. 1.
    For every $\mathcal{E}$ locally free coherent $\mathcal{O}_{X}$-module of rank $r+1$, we have $s_{i}\left(\mathcal{E}\right)=0$ for $i<0$ and $s_{0}\left(\mathcal{E}\right)=\mathrm{id}$;
  2. 2.
    For every $\mathcal{E},\mathcal{F}$ locally free coherent $\mathcal{O}_{X}$-modules, we have
    $s_{i}\left(\mathcal{E}\right)\circ s_{j}\left(\mathcal{F}\right)=s_{j}\left(\mathcal{F}\right)\circ s_{i}\left(\mathcal{E}\right)$
    for all $i,j\geq-r$;
  3. 3.
    If $Y$ is a projective variety and $f:Y\to X$ is a morphism, then, for every $\mathcal{E}$ locally free coherent $\mathcal{O}_{X}$-module and every cycle $\alpha$ on $X$, then Projection Formula holds
    $f_{\ast}\left(s_{i}\left(f^{\ast}\mathcal{E}\right)\left(\alpha\right)\right)=s_{i}\left(\mathcal{E}\right)\left(f_{\ast}\alpha\right);$
  4. 4.
    For every line bundle $\mathcal{L}$ on $X$ and $i\geq 0$, we have
    $s_{i}\left(\mathcal{L}\right):A_{k}\left(X\right)$ $\to A_{k-i}\left(X\right)$
    $\alpha$ $\mapsto\left(-1\right)^{i}iD\cdot\alpha.$
Proof.
See [Ful, Proposition 3.1, pag. 48]. ∎
Corollary 3.15.
Let $X$ be a projective variety, $\mathcal{E}$ a locally free coherent $\mathcal{O}_{X}$-module of rank $r+1$ and $\pi:\mathbb{P}_{X}\left(\mathcal{E}\right)\to X$ the associated projective bundle. Then
  1. 1.
    $\pi_{\ast}:A_{\bullet}\left(\mathbb{P}_{X}\left(\mathcal{E}\right)\right)\to A_{\bullet}\left(X\right)$ is surjective;
  2. 2.
    $\pi^{\ast}:A_{\bullet}\left(X\right)\to A_{\bullet}\left(\mathbb{P}_{X}\left(\mathcal{E}\right)\right)$ is split injective.
Proof.
It follows from Proposition 3.14 that for all $\alpha\in A_{\bullet}\left(X\right)$
$\alpha=\pi_{\ast}\left(rD_{\mathcal{E}}\cdot\pi^{\ast}\alpha\right),$
hence $\pi_{\ast}$ is surjective. Let
$F:A_{k+r}\left(\mathbb{P}_{X}\left(\mathcal{E}\right)\right)$ $\to A_{k}\left(X\right)$
$\beta$ $\mapsto\pi_{\ast}\left(rD_{\mathcal{E}}\cdot\beta\right),$
then $F\circ\pi^{\ast}=\mathrm{id}$. ∎
Theorem 3.16 ((Splitting construction)).
Let $X$ be a projective variety and let $\mathcal{E}$ be a locally free coherent $\mathcal{O}_{X}$-module of rank $r+1$. Then there exists a projective variety $Y$ and a morphism $f:Y\to X$ such that
  1. 1.
    $f$ decomposes as a sequence of projective bundles;
  2. 2.
    the map $f_{\ast}:A_{\bullet}\left(Y\right)\to A_{\bullet}\left(X\right)$ is surjective;
  3. 3.
    the map $f^{\ast}:A_{\bullet}\left(X\right)\to A_{\bullet}\left(Y\right)$ is split injective;
  4. 4.
    there exists a sequence
    $0=\mathcal{E}_{0}\subseteq\mathcal{E}_{1}\subseteq\dots\subseteq\mathcal{E}_{r}=f^{\ast}\mathcal{E}$
    of locally free $\mathcal{O}_{Y}$-modules such that $\frac{\mathcal{E}_{i}}{\mathcal{E}_{i-1}}$ is locally free of rank 1 for every $i=1,\dots,r$.
Proof.
We work by induction on $r$. If $r=0$, then there is nothing to prove. Suppose $r>0$ and set
$Y^{\prime}=\mathbb{P}_{X}\left(\mathcal{E}^{\lor}\right)\overset{f^{\prime}}{\longrightarrow}X.$
By the above, we have an exact sequence of locally free sheaves [^13] Exercise: why is $\tilde{\mathcal{E}}$ also locally free?
$0\to\tilde{\mathcal{E}}\to\left(f^{\prime}\right)^{\ast}\mathcal{E}^{\lor}\to\mathcal{O}_{Y^{\prime}}\left(1\right)\to 0,$
moreover, since $\mathrm{rk}\mathcal{O}_{Y^{\prime}}\left(1\right)=1$, we have
$\mathrm{rk}\tilde{\mathcal{E}}$ $=\mathrm{rk}\left(f^{\prime}\right)^{\ast}\mathcal{E}-1$
$=\mathrm{rk}\mathcal{E}-1$
$=r,$
we conclude by induction and Corollary 3.15. ∎
3.3 Chern and Todd classes
Definition 3.17.
Let $X$ be a projective variety and let $\mathcal{E}$ be a locally free coherent $\mathcal{O}_{X}$-module of rank $r+1$. Consider the Segre classes of $\mathcal{E}$ as endomorphisms of $A_{\bullet}\left(X\right)$, we define recursively
$c_{0}\left(\mathcal{E}\right)$ $=\mathrm{id}$
$c_{1}\left(\mathcal{E}\right)$ $=-s_{1}\left(\mathcal{E}\right)$
$c_{2}\left(\mathcal{E}\right)$ $=-s_{2}\left(\mathcal{E}\right)+s_{1}\left(\mathcal{E}\right)^{2}$
$c_{n}\left(\mathcal{E}\right)$ $=-\sum_{i=1}^{n}s_{i}\left(\mathcal{E}\right)c_{n-i}\left(\mathcal{E}\right),$
called the Chern classes of $\mathcal{E}$.
Remark 3.18.
We see that $c_{i}\left(\mathcal{E}\right)$ gives a map $A_{k}\left(X\right)\to A_{k-i}\left(X\right)$, therefore we have the feeling that the sequence of $c_{i}\left(\mathcal{E}\right)$ is bounded: it is true (namely, one can show that $c_{i}\left(\mathcal{E}\right)=0$ for $i>\mathrm{rk}\mathcal{E}$, see [Ful, Theorem 3.2, pag. 50]). Therefore the total Chern class
$c\left(\mathcal{E}\right)=\sum_{i\geq 0}c_{i}\left(\mathcal{E}\right)$
is well defined.
The following properties are just a formal translation of the corresponding ones in Proposition 3.14
Proposition 3.19.
Let $X$ be a projective variety:
  1. 1.
    For every $\mathcal{E},\mathcal{F}$ locally free coherent $\mathcal{O}_{X}$-modules, we have
    $c_{i}\left(\mathcal{E}\right)\circ c_{j}\left(\mathcal{F}\right)=c_{j}\left(\mathcal{F}\right)\circ c_{i}\left(\mathcal{E}\right)$
    for all $i,j$;
  2. 2.
    If $Y$ is a projective variety and $f:Y\to X$ is a morphism, then, for every $\mathcal{E}$ locally free coherent $\mathcal{O}_{X}$-module and every cycle $\alpha$ on $X$, then Projection Formula holds
    $f_{\ast}\left(c_{i}\left(f^{\ast}\mathcal{E}\right)\left(\alpha\right)\right)=c_{i}\left(\mathcal{E}\right)\left(f_{\ast}\alpha\right).$
Example 3.20.
Let $\mathcal{L}$ be an invertible sheaf on a smooth projective variety $X$, then $\mathcal{L}\simeq\mathcal{O}_{X}\left(D\right)$ for some divisor $D$. In view of Remark 3.18, we see that $c_{i}\left(\mathcal{L}\right)=0$ for $i\geq 2$, therefore
$c\left(\mathcal{L}\right)$ $=1+c_{1}\left(\mathcal{L}\right)$
$=1-s_{1}\left(\mathcal{L}\right)$
where $s_{1}\left(\mathcal{L}\right)$ is just the intersection with $D$.
Notation.
By abuse of notation, whenever there exists a divisor $D$ such that
$c_{i}\left(\mathcal{E}\right):A_{\bullet}\left(X\right)$ $\to A_{\bullet}\left(X\right)$
$\alpha$ $\mapsto D\cdot\alpha,$
we’ll write $c_{i}\left(\mathcal{E}\right)=D$. With this notation, in view of Example 3.20, $c\left(\mathcal{O}_{X}\left(D\right)\right)=1+D$ for any divisor $D$.
The main property of the total Chern class is
Proposition 3.21 ((Whitney sum)).
Let
$0\to\mathcal{F}\to\mathcal{G}\to\mathcal{H}\to 0$
be an exact sequence of locally free coherent $\mathcal{O}_{X}$-modules on a smooth projective variety $X$. Then
$c\left(\mathcal{G}\right)=c\left(\mathcal{F}\right)\circ c\left(\mathcal{H}\right).$
Proof.
It is [Ful, Theorem 3.2 (e), pag. 51] for $t=1$. ∎
Example 3.22.
Let $X=\mathbb{P}^{n}$ and let us compute $c\left(\mathcal{T}_{X}\right)$, where $\mathcal{T}$ denotes the tangent sheaf. We have the Euler exact sequence
$0\to\mathcal{O}_{X}\to\mathcal{O}_{X}\left(1\right)^{\oplus n+1}\to\mathcal{T}_{X}\to 0$
and hence, in view of Proposition 3.21, we have
$c\left(\mathcal{O}_{X}\left(1\right)^{\oplus n+1}\right)=c\left(\mathcal{O}_{X}\right)\circ c\left(\mathcal{T}_{X}\right),$
with $c\left(\mathcal{O}_{X}\right)=1$ and
$c\left(\mathcal{O}_{X}\left(1\right)^{\oplus n+1}\right)$ $=c\left(\mathcal{O}_{X}\left(1\right)^{\oplus n+1}\right)=\left(1+H\right)^{n+1}.$
Then
$c_{k}\left(\mathcal{T}_{X}\right)=\binom{n+1}{k}\cdot H^{k},$
where $H^{k}$ is the class of a $k$-dimensional linear subspace [^14] Here we’re cheating a bit: we haven’t proved that all $k$-dimensional linear subspaces of $\mathbb{P}^{n}$ are equivalent. Take it as an act of faith or see [Ful, Example 1.9.3, pag. 23]..
Remark 3.23.
The way one uses Proposition 3.21 and Theorem 3.16 is as follows: in view of Example 3.20, given a locally free coherent $\mathcal{O}_{X}$-module $\mathcal{E}$ of rank $r+1$ on a smooth projective variety $X$, we have a product
$c\left(\mathcal{E}\right)=\prod_{i=1}^{r+1}\left(1+\alpha_{i}\right)$
on some projective bundle over $X$. In view of Corollary 3.15, we see that the $k$-th symmetric polynomial in the $\alpha_{i}$’s (treating them as formal variables) is actually $c_{k}\left(\mathcal{E}\right)$, hence every symmetric polynomial in the $\alpha_{i}$’s in expressible in terms of the Chern classes. For a (more) formal explanation of this principle one can see [Ful, Remark 3.2.3, pag. 54]. One can also observe that, using this trick, given a morphism $f:X\to Y$ between projective varieties and a locally free coherent $\mathcal{O}_{Y}$-module $\mathcal{E}$, we have
$c_{k}\left(f^{\ast}\mathcal{E}\right)=f^{\ast}c_{k}\left(\mathcal{E}\right).$
A proof of this can be found in [EH, Section 5.4].
Example 3.24.
Using the splitting principle one can see how Chern roots behave under linear algebra operations: if $\mathcal{E}$ has Chern roots $\alpha_{1},\dots,\alpha_{n}$, then
  • $\mathcal{E}^{\lor}$ has Chern roots $\alpha_{1},\dots,-\alpha_{n}$;
  • $S^{k}\mathcal{E}$ (the symmetric power) has Chern roots $\alpha_{i_{1}}+\dots+\alpha_{i_{k}}$ for $1\leq i_{1}\leq\dots\leq i_{k}\leq n$;
  • $\bigwedge^{k}\mathcal{E}$ (the exterior power) has Chern roots $\alpha_{i_{1}}+\dots+\alpha_{i_{k}}$ for $1\leq i_{1}<\dots<i_{k}\leq n$.
For a proof of this formulas see [Ful, Remark 3.2.3, pag. 54]
Definition 3.25.
The formal variables $\alpha_{1},\dots,\alpha_{r+1}$ introduced in Remark 3.23 are called the Chern roots of $\mathcal{E}$. We set [^15] With this notation we mean $\exp\left(x\right)$ $=1+x+\frac{1}{2}x^{2}+\dots$ $\frac{x}{1-\exp\left(-x\right)}$ $=1+\frac{1}{2}x+\frac{1}{12}x^{2}+\dots$ as formal power series in $\mathbb{Q}\left[\left[x\right]\right]$.
$\mathrm{ch}\left(\mathcal{E}\right)$ $=\sum_{i=1}^{r+1}\exp\left(\alpha_{i}\right)$
$\mathrm{td}\left(\mathcal{E}\right)$ $=\prod_{i=1}^{r+1}\frac{\alpha_{i}}{1-\exp\left(-\alpha_{i}\right)}$
as endomorphisms of $A_{\bullet}\left(X\right)\otimes\mathbb{Q}$. Ther are called the Chern and Todd character, respectively.
Remark 3.26.
The Chern and Todd characters are actually given by finite sums, as follows from Remark 3.18. Moreover they are clearly symmetric in the $\alpha_{i}$’s, so that they are expressible using the Chern characters.
Proposition 3.27.
Let
$0\to\mathcal{F}\to\mathcal{G}\to\mathcal{H}\to 0$
be an exact sequence of locally free coherent $\mathcal{O}_{X}$-modules on a smooth projective variety $X$. Then
$\mathrm{ch}\left(\mathcal{G}\right)$ $=\mathrm{ch}\left(\mathcal{F}\right)+\mathrm{ch}\left(\mathcal{H}\right)$
$\mathrm{td}\left(\mathcal{G}\right)$ $=\mathrm{td}\left(\mathcal{F}\right)\circ\mathrm{td}\left(\mathcal{H}\right).$
Proof.
Applying the splitting construction three times, we find a morphism $f:Y\to X$ from a projective variety $Y$ over which $\mathcal{F},\mathcal{G}$ and $\mathcal{H}$ have a filtration like in Theorem 3.16. There are (at least) two good reasons for which our starting exact sequence stays exact when pulled-back to $Y$: first note that $f$ is constructed as a sequence of projective bundles, therefore, locally, it looks like
$\mathbb{P}_{A}^{r}\to\mathrm{Spec}\left(A\right)$
and hence it is flat. Otherwise one notes that we’re working with locally free modules, therefore our sequence is locally split. Since the functor $f^{\ast}$ is additive, it preserve the exactness of locally split exact sequences. [^16] Recall that, when working with $\mathcal{O}_{Y}$-modules, exactness is checked locally. In view of Proposition 3.21, we deduce that the set of Chern roots of $\mathcal{G}$ is just the union of those of $\mathcal{F}$ and those of $\mathcal{H}$, therefore the Proposition reduces to a formal rearrangement in the definitions of $\mathrm{ch}\left(\mathcal{G}\right)$ and $\mathrm{td}\left(\mathcal{G}\right)$. ∎
Black Box: turning the tables So far we have defined Chern classes as endomorphisms of $A_{\bullet}\left(X\right)$. To have a more geometric flavour, we shall (and need to) see them as actual elements of $A_{\bullet}\left(X\right)$, and we can address this as follows: for a smooth projective variety $X$ one would like to put a commutative ring structure on $A^{\bullet}\left(X\right)$ taking, for $D,E$ irreducible cycles of codimension $d$ and $e$ respectively, $\left[D\right]\cdot\left[E\right]=\sum n_{i}\left[W_{i}\right]$ where the $W_{i}$’s are the irreducible components of $D\cap E$ and the $n_{i}$’s are integers defined in some smart way, called intersection multiplicities. This turns out to be possible and paves the way to a very deep subject, namely the so-called intersection theory. Anyway, in our specific case (that is, smooth projective varieties over an algebraically closed field) the theory goes without any great harm. One may want to page through [Ful, Chapters 6,7] for the whole geometric picture, while in the book [Ser] an algebraic definition of intersection multiplicity, reminiscent of the ring structure introduced in Proposition 2.12, is given and generalized, leading to Serre’s Intersection Multiplicity Conjectures, which is at present one of the most outstanding open problem in Commutative and Homological Algebra (see also Section 8 in [Hoc] for a discussion of the problem and the state of the art). Since $A_{\bullet}\left(X\right)$ and $A^{\bullet}\left(X\right)$ are the same abelian group, this gives a ring structure on $A_{\bullet}\left(X\right)$ where $\left[X\right]$ is the unit element. We will identify $c_{i}\left(\mathcal{E}\right)$ with $c_{i}\left(\mathcal{E}\right)\cdot\left[X\right]\in A_{\bullet}\left(X\right),$ then all the statements and properties we have seen in this section carry over as statements about cycles.
Remark 3.28.
In view of Proposition 3.27, the previous trick allows us to define the Chern character $\mathrm{ch}\left(\bullet\right)$ as a group homomorphism
$\mathrm{ch}\left(\bullet\right):K_{0}\left(X\right)\to A_{\bullet}\left(X\right)\otimes\mathbb{Q}.$
Moreover, to a smooth projective variety $X$, we can associate an element
$\mathrm{td}\left(X\right)=\mathrm{td}\left(\mathcal{T}_{X}\right)\in A_{\bullet}\left(X\right)\otimes\mathbb{Q}$
called the Todd class of $X$, where $\mathcal{T}_{X}$ is the tangent sheaf of $X$.
Example 3.29.
It follows from Example 3.22 that the Chern roots of $\mathcal{T}_{\mathbb{P}^{n}}$ are just $n+1$ times the class of a hyperplane $H$, therefore
$\mathrm{td}\left(\mathbb{P}^{n}\right)=\frac{H^{n+1}}{\left(1-\exp\left(-H\right)\right)^{n+1}},$
which, as usual, has to be meant as a Taylor expansion in $A_{\bullet}\left(\mathbb{P}^{n}\right)\otimes\mathbb{Q}$.
4 Hirzebruch-Riemann-Roch for smooth projective varieties
The aim of this final section is to state and prove the Hirzebruch-Riemann-Roch theorem for smooth projective varieties. This is stated as
Theorem 4.1.
Let $X$ be a smooth projective variety, then the diagram
[…]
is commutative, that is, for every locally free coherent $\mathcal{O}_{X}$-module $\mathcal{E}$, we have
$\chi\left(X,\mathcal{E}\right)=\int_{X}\mathrm{ch}\left(\mathcal{E}\right)\cdot\mathrm{td}\left(X\right).$
4.1 The case of $\mathbb{P}^{n}$
Theorem 4.2.
The diagram
${K_{0}\left(\mathbb{P}^{n}\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\chi\left(\mathbb{P}^{n},\bullet\right)}$$\scriptstyle{\mathrm{ch}\left(\bullet\right)\cdot\mathrm{td}\left(\mathbb{P}^{n}\right)}$${A_{\bullet}\left(\mathbb{P}^{n}\right)\otimes\mathbb{Q}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\int_{\mathbb{P}^{n}}}\textstyle$$\mathbb{Z}\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${\mathbb{Q}}$
is commutative.
Proof.
Let $\mathcal{F}$ be a locally free coherent $\mathcal{O}_{\mathbb{P}^{n}}$-module, we want to show that
$\chi\left(\mathbb{P}^{n},\mathcal{F}\right)=\int_{\mathbb{P}^{n}}\mathrm{ch}\left(\mathcal{F}\right)\cdot\mathrm{td}\left(\mathbb{P}^{n}\right).$
In view of Theorem 2.18 and the fact that $\mathrm{ch}$ is a group homomorphism, it is enough to show it for $\mathcal{F}=\mathcal{O}_{\mathbb{P}^{n}}\left(d\right)$ with $d=0,\dots,n$. On the one hand we know from Lemma 2.17 that
$\chi\left(\mathbb{P}^{n},\mathcal{O}_{\mathbb{P}^{n}}\left(d\right)\right)=\binom{n+d}{n}$,
on the other hand $\mathrm{ch}\left(\mathcal{O}_{\mathbb{P}^{n}}\left(d\right)\right)=\exp\left(dH\right)$ where $H\subseteq\mathbb{P}^{n}$ is a hyperplane. In view of Example 3.29 we have
$\mathrm{td}\left(\mathbb{P}^{n}\right)=\frac{H^{n+1}}{\left(1-\exp\left(-H\right)\right)^{n+1}},$
and hence [^17] To show the equality $\int_{\mathbb{P}^{n}}\mathrm{ch}\left(\mathcal{O}_{\mathbb{P}^{n}}\left(d\right)\right)\cdot\mathrm{td}\left(\mathbb{P}^{n}\right)=\binom{n+d}{n}$ one uses the following trick: the point is to compute the $n$-th coefficient in the Taylor expansion of the function $f\left(x\right)=e^{dx}\dfrac{x^{n+1}}{\left(1-e^{-x}\right)^{n+1}}$, but this is exactly the residue $\mathrm{res}_{0}\left(\dfrac{e^{dx}}{\left(1-e^{-x}\right)^{n+1}}\right)$, which can be computed by the change of variables $y=1-e^{-x}$.
$\int_{\mathbb{P}^{n}}\mathrm{ch}\left(\mathcal{O}_{\mathbb{P}^{n}}\left(d\right)\right)\cdot\mathrm{td}\left(\mathbb{P}^{n}\right)=\binom{n+d}{n}=\chi\left(\mathbb{P}^{n},\mathcal{O}_{\mathbb{P}^{n}}\left(d\right)\right).$
4.2 The case of smooth projective varieties
Lemma 4.3.
Let $X$ be a smooth projective variety and let $j:X\to\mathbb{P}^{n}$ be a closed embedding. If the diagram
${K_{0}\left(X\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{ch}\left(\bullet\right)\cdot\mathrm{td}\left(X\right)}$$\scriptstyle{j_{!}}$${A_{\bullet}\left(X\right)\otimes\mathbb{Q}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j_{\ast}\otimes\mathbb{Q}}$${K_{0}\left(\mathbb{P}^{n}\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{ch}\left(\bullet\right)\cdot\mathrm{td}\left(\mathbb{P}^{n}\right)}$${A_{\bullet}\left(\mathbb{P}^{n}\right)\otimes\mathbb{Q}}$
is commutative, then for every locally free coherent $\mathcal{O}_{X}$-module $\mathcal{E}$, we have
$\chi\left(X,\mathcal{E}\right)=\int_{X}\mathrm{ch}\left(\mathcal{E}\right)\cdot\mathrm{td}\left(X\right).$
Proof.
We see, for example from Proposition 2.16 and [Har, Lemma III.2.10, pag. 209], that the composition
${K_{0}\left(X\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j_{!}}$${K_{0}\left(\mathbb{P}^{n}\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\chi\left(\mathbb{P}^{n},\bullet\right)}\textstyle{\mathbb{Z}}$
is just $\chi\left(X,\bullet\right)$. We are then reduced to showing that, for $\alpha\in A_{\bullet}\left(X\right)$, we have
$\int_{Y}j_{\ast}\left(\alpha\right)=\int_{X}\alpha,$
but, in view of the very definition of $\int$, this is Lemma 3.5. ∎
Lemma 4.4.
Let $j:Y\to X$ be a closed embedding between smooth projective varieties. If, for every locally free coherent $\mathcal{O}_{Y}$-module $\mathcal{E}$, there exists a locally free coherent resolution
$0\to\mathcal{F}_{r}\to\dots\to\mathcal{F}_{1}\to\mathcal{F}_{0}\to j_{\ast}\mathcal{E}\to 0$
by $\mathcal{O}_{X}$-modules such that
$\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k}\right)=j_{\ast}\left(\frac{\mathrm{ch}\left(\mathcal{E}\right)}{\mathrm{td}\left(\mathcal{N}_{X|Y}\right)}\right),$
then the diagram
${K_{0}\left(Y\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{ch}\left(\bullet\right)\cdot\mathrm{td}\left(Y\right)}$$\scriptstyle{j_{!}}$${A_{\bullet}\left(Y\right)\otimes\mathbb{Q}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j_{\ast}\otimes\mathbb{Q}}$${K_{0}\left(X\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{ch}\left(\bullet\right)\cdot\mathrm{td}\left(X\right)}$${A_{\bullet}\left(X\right)\otimes\mathbb{Q}}$
is commutative.
Proof.
We have an exact sequence
$0\to\mathcal{T}_{Y}\to j^{\ast}\mathcal{T}_{X}\to\mathcal{N}_{Y|X}\to 0$
of locally free $\mathcal{O}_{Y}$-modules, therefore, in view of Proposition 3.27 and Remark 3.23 we see that
$j^{\ast}\mathrm{td}\left(X\right)=\mathrm{td}\left(j^{\ast}\mathcal{T}_{X}\right)=\mathrm{td}\left(Y\right)\cdot\mathrm{td}\left(\mathcal{N}_{Y|X}\right).$ (1)
We need to see that, for $y\in K_{0}\left(Y\right)$, the equality
$j_{\ast}\left(\mathrm{ch}\left(y\right)\cdot\mathrm{td}\left(Y\right)\right)=\mathrm{ch}\left(j_{!}\left(y\right)\right)\cdot\mathrm{td}\left(X\right)$
holds. In view of Proposition 2.9, we only need to check it when $y$ is the class of a locally free coherent $\mathcal{O}_{Y}$-module $\mathcal{E}$. Let
$0\to\mathcal{F}_{r}\to\dots\to\mathcal{F}_{1}\to\mathcal{F}_{0}\to j_{\ast}\mathcal{E}\to 0$
be a resolution as in the statement, then we have
$j_{!}\left[\mathcal{E}\right]=\sum_{k=0}^{r}\left(-1\right)^{k}\left[\mathcal{F}_{k}\right].$
In view of Proposition 3.27 we have
$\mathrm{ch}\left(j_{!}\left[\mathcal{E}\right]\right)\cdot\mathrm{td}\left(X\right)$ $=\mathrm{ch}\left(\sum_{k=0}^{r}\left(-1\right)^{k}\left[\mathcal{F}_{k}\right]\right)\cdot\mathrm{td}\left(X\right)$
$=\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k}\right)\cdot\mathrm{td}\left(X\right)$
$=j_{\ast}\left(\mathrm{ch}\left(\mathcal{E}\right)\cdot\mathrm{td}\left(\mathcal{N}_{X|Y}\right)^{-1}\right)\cdot\mathrm{td}\left(X\right)$
$\overset{\ast}{=}j_{\ast}\left(\mathrm{ch}\left(\mathcal{E}\right)\cdot\mathrm{td}\left(\mathcal{N}_{X|Y}\right)^{-1}\cdot j^{\ast}\mathrm{td}\left(X\right)\right)$
$=j_{\ast}\left(\mathrm{ch}\left(\mathcal{E}\right)\cdot\mathrm{td}\left(Y\right)\right)$
where the equality marked with $\ast$ follows from Projection Formula in Proposition 3.19 and the last one is Formula (1). ∎
We need an intermediate step
Lemma 4.5.
Let $X$ be a smooth projective variety and let $\mathcal{E}$ be a locally free coherent sheaf of rank $r$, set $Y=\mathbb{P}_{X}\left(\mathcal{E}\oplus\mathcal{O}_{X}\right)$. Then the diagram
${K_{0}\left(X\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{ch}\left(\bullet\right)\cdot\mathrm{td}\left(X\right)}$$\scriptstyle{i_{!}}$${A_{\bullet}\left(X\right)\otimes\mathbb{Q}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i_{\ast}\otimes\mathbb{Q}}$${K_{0}\left(Y\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{ch}\left(\bullet\right)\cdot\mathrm{td}\left(Y\right)}$${A_{\bullet}\left(Y\right)\otimes\mathbb{Q}}$
is commutative, where the map $i:X\to Y$ is the one induced by the inclusion $\mathcal{O}_{X}\subseteq\mathcal{E}\oplus\mathcal{O}_{X}$.
Proof.
Note that $i$ is a section for the structure map $\pi:Y\to X$, moreover from Remark 3.12 we have an exact sequence of locally free sheaves
$0\to\mathcal{L}\to\pi^{\ast}\left(\mathcal{E}\oplus\mathcal{O}_{X}\right)\to\mathcal{O}_{Y}\left(1\right)\to 0$
which, dualized, gives181818
Exercise.
Show that $\pi^{\ast}\left(\mathcal{E}^{\lor}\right)\simeq\left(\pi^{\ast}\mathcal{E}\right)^{\lor}$.
$0\to\mathcal{O}_{Y}\left(-1\right)\to\pi^{\ast}\left(\mathcal{E}^{\lor}\oplus\mathcal{O}_{X}\right)\to\mathcal{L}^{\lor}\to 0.$
Note that $\pi^{\ast}\left(\mathcal{E}\oplus\mathcal{O}_{X}\right)$ has a canonical section
$\mathcal{O}_{Y}=\pi^{\ast}\mathcal{O}_{X}\to\pi^{\ast}\left(\mathcal{E}^{\lor}\oplus\mathcal{O}_{X}\right).$
[…]
$0\to\bigwedge^{r}\mathcal{L}\to\dots\to\bigwedge^{2}\mathcal{L}\to\mathcal{L}\overset{s}{\longrightarrow}\mathcal{O}_{Y}\to\mathcal{O}_{Z\left(s\right)}\to 0.$
Let $\mathcal{F}$ be a locally free coherent $\mathcal{O}_{X}$-module, then
$\pi^{\ast}\mathcal{F}\otimes\mathcal{O}_{Z\left(s\right)}$ $\simeq\pi^{\ast}\mathcal{F}\otimes i_{\ast}\mathcal{O}_{X}$
$\simeq i_{\ast}\left(i^{\ast}\pi^{\ast}\mathcal{F}\otimes\mathcal{O}_{X}\right)$
$\simeq i_{\ast}\mathcal{F}$
and hence we get a locally free resolution on $Y$
$0\to\pi^{\ast}\mathcal{F}\otimes\bigwedge^{r}\mathcal{L}\to\dots\to\pi^{\ast}\mathcal{F}\otimes\bigwedge^{2}\mathcal{L}\to\pi^{\ast}\mathcal{F}\otimes\mathcal{L}\overset{s}{\longrightarrow}\pi^{\ast}\mathcal{F}\to i_{\ast}\mathcal{F}\to 0.$
In order to apply Lemma 4.4 we need to see that
$\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\pi^{\ast}\mathcal{F}\otimes\bigwedge^{k}\mathcal{L}\right)=i_{\ast}\left(\frac{\mathrm{ch}\left(\mathcal{F}\right)}{\mathrm{td}\left(\mathcal{N}_{Y|X}\right)}\right).$
First we show that $\mathcal{N}_{Y|X}\simeq i^{\ast}\mathcal{L}^{\lor}$: we have an exact sequence
$0\to\mathcal{I}\to\mathcal{O}_{Y}\to\mathcal{O}_{X}\to 0$
on $Y$, giving that
$\mathrm{Tor}_{1}^{\mathcal{O}_{Y}}\left(\mathcal{O}_{X},\mathcal{O}_{X}\right)\simeq\frac{\mathcal{I}}{\mathcal{I}^{2}}=\mathcal{N}_{Y|X}^{\lor},$
on the other hand, in view of Remark A.3, we have
$i^{\ast}\mathcal{L}=\mathrm{Tor}_{1}^{\mathcal{O}_{Y}}\left(\mathcal{O}_{X},\mathcal{O}_{X}\right)$
and hence $\mathcal{N}_{Y|X}\simeq i^{\ast}\mathcal{L}^{\lor}$. One can show that
$i_{\ast}\left(\frac{\mathrm{ch}\left(\mathcal{F}\right)}{\mathrm{td}\left(i^{\ast}\mathcal{L}^{\lor}\right)}\right)=\frac{\mathrm{ch}\left(\pi^{\ast}\mathcal{F}\right)}{\mathrm{td}\left(\mathcal{L}^{\lor}\right)}\cdot c_{r}\left(\mathcal{L}^{\lor}\right)$
(cf. [Ful, Section 14.1, pag. 244]), so that we need to show the equality
$\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\bigwedge^{k}\mathcal{L}\right)=\frac{c_{r}\left(\mathcal{L}^{\lor}\right)}{\mathrm{td}\left(\mathcal{L}^{\lor}\right)},$
but, if $a_{1},\dots,a_{r}$ are the Chern roots of $\mathcal{L}$, then, from Example 3.24,
$\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\bigwedge^{k}\mathcal{L}\right)$ $=\sum_{k=0}^{r}\left(-1\right)^{k}\left(\sum_{1\leq i_{1}<\dots<i_{k}\leq n}\exp\left(a_{i_{1}}+\dots+a_{i_{k}}\right)\right)$
$=\prod_{i=1}^{r}\left(1-\exp\left(a_{i}\right)\right)$
$=a_{1}\cdot\dots\cdot a_{r}\cdot\prod_{i=1}^{r}\frac{1-\exp\left(a_{i}\right)}{a_{i}}$
$=c_{r}\left(\mathcal{L}^{\lor}\right)\cdot\mathrm{td}\left(\mathcal{L}^{\lor}\right)^{-1}.$
Theorem 4.6.
Let $X$ be a smooth projective variety, then the diagram
[…]
is commutative, that is, for every locally free coherent $\mathcal{O}_{X}$-module $\mathcal{E}$, we have
$\chi\left(X,\mathcal{E}\right)=\int_{X}\mathrm{ch}\left(\mathcal{E}\right)\cdot\mathrm{td}\left(X\right).$
Proof.
Let $j:X\to\mathbb{P}^{n}$ be a closed embedding, $\mathcal{E}$ be a locally free coherent $\mathcal{O}_{X}$-module. Fix a point $0\in\mathbb{P}^{1}$ and consider the blowing up
$\pi:M=\mathrm{Bl}_{X\times\left\{0\right\}}\left(\mathbb{P}^{n}\times\mathbb{P}^{1}\right)\to\mathbb{P}^{n}\times\mathbb{P}^{1}$
together with the projection map $q:M\to\mathbb{P}^{1}$. For $P\neq 0$ we have $q^{-1}\left(P\right)\simeq\mathbb{P}^{n}$, while $q^{-1}\left(0\right)$ is the union of the exception divisor $E\simeq\mathbb{P}_{X}\left(\mathcal{N}_{X|\mathbb{P}^{n}}\oplus\mathcal{O}_{X}\right)$ and $Y\simeq\mathrm{Bl}_{X}\left(\mathbb{P}^{n}\right)$. Note that the product $j\times\mathbb{P}^{1}:X\times\mathbb{P}^{1}\to\mathbb{P}^{n}\times\mathbb{P}^{1}$ makes the preimage of $X\times\left\{0\right\}$ locally principal, hence there exists a unique $v:X\times\mathbb{P}^{1}\to M$ making the diagram
[…]
commute. It is moreover clear that over $P\neq 0\in\mathbb{P}^{1}$ the map $v_{P}$ gives the embedding $X\times\left\{P\right\}\to\mathbb{P}^{n}\times\left\{P\right\}$ induced by $j$, while $v_{0}$ is the embedding
$X\times\left\{0\right\}\simeq X\to\mathbb{P}_{X}\left(\mathcal{N}_{X|\mathbb{P}^{n}}\oplus\mathcal{O}_{X}\right)$
considered in Lemma 4.5. Denote with $p:X\times\mathbb{P}^{1}\to X$ the projection, then, since $M$ is smooth projective, there exists a locally free resolution by coherent $\mathcal{O}_{M}$-modules
$0\to\mathcal{F}_{r}\to\dots\to\mathcal{F}_{0}\to v_{\ast}p^{\ast}\mathcal{E}\to 0$
[…]
$\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k}\right)\cdot q^{\ast}\left(\left[0\right]-\left[\infty\right]\right)=0$
in $A_{\bullet}\left(M\right)\otimes\mathbb{Q}$. As we noted, we have
$q^{\ast}\left[0\right]$ $=\left[E\right]+\left[Y\right]$
$q^{\ast}\left[\infty\right]$ $=\left[\mathbb{P}^{n}\right],$
hence we have a relation
$\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|Y}\right)\cdot\left[Y\right]+\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|E}\right)\cdot\left[E\right]=\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|\mathbb{P}^{n}}\right)\cdot\left[\mathbb{P}^{n}\right].$
Note that the sheaf $v_{\ast}p^{\ast}\mathcal{E}$ is zero on $Y$, therefore
$0\to\mathcal{F}_{r|Y}\to\dots\to\mathcal{F}_{0|Y}\to 0$
is exact and hence $\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|Y}\right)\cdot\left[Y\right]=0$. In view of Lemma 4.5 we have
$\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|E}\right)\cdot\left[E\right]=\frac{\mathrm{ch}\left(\mathcal{E}\right)}{\mathrm{td}\left(\mathcal{N}_{X|\mathbb{P}^{n}}\right)}\cdot\left[X\right],$
therefore
$\frac{\mathrm{ch}\left(\mathcal{E}\right)}{\mathrm{td}\left(\mathcal{N}_{X|\mathbb{P}^{n}}\right)}\cdot\left[X\right]=\sum_{k=0}^{r}\left(-1\right)^{k}\mathrm{ch}\left(\mathcal{F}_{k|\mathbb{P}^{n}}\right)\cdot\left[\mathbb{P}^{n}\right]$
which, pushed forward along $g:M\to\mathbb{P}^{n}$, gives the relation as in the hypothesis of Lemma 4.4. ∎
4.3 Some consequences
4.3.1 Explicit formulas in low dimension
Theorem 4.7 ((Riemann-Roch for curves)).
Let $\mathcal{L}\in\mathrm{Pic}\left(C\right)$ be locally free sheaf on $C$ of rank 1, then
$\chi\left(C,\mathcal{L}\right)=\deg\mathcal{L}+\chi\left(C,\mathcal{O}_{C}\right).$
Proof.
We only need to see that
$\int_{C}\mathrm{ch}\left(\mathcal{L}\right)\cdot\mathrm{td}\left(\mathcal{C}\right)=\deg\mathcal{L}+\chi\left(C,\mathcal{O}_{C}\right).$
Let $D$ be a divisor such that $\mathcal{L}\simeq\mathcal{O}_{C}\left(D\right)$, then
$\mathrm{ch}\left(\mathcal{L}\right)=1+c_{1}\left(\mathcal{L}\right)$
and
$\mathrm{td}\left(C\right)=1+\frac{1}{2}c_{1}\left(\mathcal{T}_{C}\right).$
Since $C$ has dimension 1, we have $A_{2}\left(C\right)=0$ and hence
$\int_{C}\mathrm{ch}\left(\mathcal{L}\right)\cdot\mathrm{td}\left(C\right)$ $=\int_{C}\left(1+c_{1}\left(\mathcal{L}\right)\right)\cdot\left(1+\frac{1}{2}c_{1}\left(\mathcal{T}_{C}\right)\right)$
$=\int_{C}\left(1+c_{1}\left(\mathcal{L}\right)+\frac{1}{2}c_{1}\left(\mathcal{T}_{C}\right)\right)$
$=\int_{C}\left(1+c_{1}\left(\mathcal{L}\right)-\frac{1}{2}c_{1}\left(\Omega_{C}^{1}\right)\right)$
$=\int_{C}\left(c_{1}\left(\mathcal{L}\right)-\frac{1}{2}c_{1}\left(\Omega_{C}^{1}\right)\right)$
$=\deg\left(D\right)-\frac{1}{2}\left(2g-2\right)$
$=\deg\left(D\right)-g+1.$
Theorem 4.8 ((Riemann-Roch for surfaces)).
Let Let $\mathcal{L}\in\mathrm{Pic}\left(S\right)$ be locally free sheaf on $S$ of rank 1, then
$\chi\left(S,\mathcal{L}\right)=\frac{1}{2}\mathcal{L}\cdot\left(\mathcal{L}\otimes_{\mathcal{O}_{S}}\omega_{S}^{\lor}\right)+\frac{1}{12}\left(K_{S}^{2}+\int_{S}c_{2}\left(\mathcal{T}_{S}\right)\right).$
In particular
$\chi\left(S,\mathcal{O}_{S}\right)=\frac{1}{12}\left(K_{S}^{2}+\int_{S}c_{2}\left(\mathcal{T}_{S}\right)\right).$
Proof.
Let $D$ be a divisor such that $\mathcal{L}\simeq\mathcal{O}_{S}\left(D\right)$, then we compute
$\int_{S}\mathrm{ch}\left(\mathcal{L}\right)\cdot\mathrm{td}\left(S\right)$ $=\int_{S}\left(1+c_{1}\left(\mathcal{L}\right)+\frac{1}{2}c_{1}\left(\mathcal{L}\right)^{2}\right)\cdot\left(1+\frac{1}{2}c_{1}\left(\mathcal{T}_{S}\right)+\frac{1}{12}\left(c_{1}\left(\mathcal{T}_{S}\right)^{2}+c_{2}\left(\mathcal{T}_{S}\right)\right)\right)$
$=\int_{S}\left(1+\frac{1}{2}c_{1}\left(\mathcal{L}\right)^{2}+\frac{1}{2}c_{1}\left(\mathcal{L}\right)\cdot c_{1}\left(\mathcal{T}_{S}\right)+\frac{1}{12}\left(c_{1}\left(\mathcal{T}_{S}\right)^{2}+c_{2}\left(\mathcal{T}_{S}\right)\right)\right)$
$=\frac{1}{2}D^{2}-\frac{1}{2}D\cdot K_{S}+\frac{1}{12}\left(K_{S}^{2}+\int_{S}c_{2}\left(\mathcal{T}_{S}\right)\right).$
The last statement follows setting $\mathcal{L}=\mathcal{O}_{S}$. ∎
Appendix A Koszul complex
Definition A.1.
Let $A$ be a commutative ring and $E$ a finitely generated $A$-module. For an element
$s\in E^{\lor}=\mathrm{Hom}_{A}\left(E,A\right)$
we define $\bigwedge^{\bullet}\left(s\right)$ as the complex191919
Exercise.
Check that this is a complex
$\dots\to\bigwedge^{n}E\overset{d_{n}}{\longrightarrow}\bigwedge^{n-1}E\to\dots\to E\overset{s}{\longrightarrow}A\to 0$
where
$d_{n}\left(e_{1}\wedge\dots\wedge e_{n}\right)=\sum_{i=1}^{n}\left(-1\right)^{i+1}s\left(e_{i}\right)e_{1}\wedge\dots\wedge\hat{e_{i}}\wedge\dots\wedge e_{n}.$
We call this complex the Koszul complex associated with $s$. We denote its homology with
$H_{k}\left(s\right)=H_{k}\left(\bigwedge^{\bullet}\left(s\right)\right)=\frac{\ker\left(d_{k}\right)}{\mathrm{Im}\left(d_{k+1}\right)}.$
We say that $s\in E^{\lor}$ is regular if $H_{k}\left(s\right)=0$ for $k\neq 0$.
Example A.2.
Let $A$ be a commutative ring and let $a_{1},\dots,a_{d}\in A$ with $\left(a_{1},\dots,a_{d}\right)\neq A$. If $s$ is associated to that sequence, we have
$H_{0}\left(s\right)=\frac{A}{\left(a_{1},\dots,a_{s}\right)}.$
Remark A.3.
Let $\bigwedge^{\bullet}\left(s\right)$ be the Koszul complex associated with the sequence $a_{1},\dots,a_{d}$ and let $I=\left(a_{1},\dots,a_{d}\right)$, then all the differentials in the complex $\bigwedge^{\bullet}\left(s\right)\otimes_{A}\frac{A}{I}$ vanish. Therefore, when $\bigwedge^{\bullet}\left(s\right)$ is exact (cfr. for example Proposition A.6), the complex $\bigwedge^{\bullet}\left(s\right)\otimes_{A}\frac{A}{I}$ computes $\mathrm{Tor}_{\bullet}^{A}\left(\frac{A}{I},\frac{A}{I}\right)$.
Note A.4.
Recall that, given a commutative ring $A$, a sequence of elements $a_{1},\dots,a_{d}$ form a regular sequence if $\left(a_{1},\dots,a_{d}\right)\neq A$ and the image of $a_{i}$ is a non-zero-divisor in
$\frac{A}{\left(a_{1},\dots,a_{i-1}\right)}.$
Remark A.5.
Let $a_{1},\dots,a_{d}\in A$ with $I=\left(a_{1},\dots,a_{d}\right)$ and set
$s:E=A^{d}$ $\to A$
$\sum c_{i}e_{i}$ $\mapsto\sum c_{i}a_{i}.$
Let now $F\subseteq E$ the free submodule generated by the first d1d-1 elements of the sequence and let $t=s_{|F}:F\to A$. Denote with $j:F\to E$ the inclusion, then it is split (since $\mathrm{coker}\left(j\right)\simeq A$) say by $g$, therefore we have a split injective map
$\wedge^{k}\left(j\right):\bigwedge^{k}F$ $\to\bigwedge^{k}E$
$f_{1}\wedge\dots\wedge f_{k}$ $\mapsto j\left(f_{1}\right)\wedge\dots\wedge j\left(f_{k}\right)$
giving a (split) monomorphism
$0\to\bigwedge^{\bullet}\left(t\right)\to\bigwedge^{\bullet}\left(s\right),$
moreover the map
$\bullet\wedge a_{d}:\bigwedge^{k-1}F\to\ker\left(\bigwedge^{k}E\overset{\wedge^{k}\left(g\right)}{\longrightarrow}\bigwedge^{k}F\right)$
can be seen to be an isomorphism and to commute with differentials, thus giving an exact sequence
$0\to\bigwedge^{\bullet}\left(t\right)\to\bigwedge^{\bullet}\left(s\right)\to\left(\bigwedge^{\bullet}\left(t\right)\right)\left[-1\right]\to 0.$ (2)
Taking homology we get an exact sequence
$H_{d}\left(t\right)\to H_{d}\left(s\right)\to H_{d-1}\left(t\right)$
where $H_{d}\left(t\right)=0$ for rank reasons, hence we have an exact sequence
$0\to H_{d}\left(s\right)\to H_{d-1}\left(t\right)\overset{\delta_{d}}{\longrightarrow}H_{d-1}\left(t\right).$ (3)
Tracking back the construction of the connecting homomorphism of the Snake Lemma one can see that
$\delta_{n}:H_{n-1}\left(t\right)$ $\to H_{n-1}\left(t\right)$
$x$ $\mapsto\pm a_{d}\cdot x,$
hence in particular $a_{d}\cdot H_{d}\left(t\right)=0$.
The connection between the Koszul complex and the properties of the sequence defining it starts with the following [^20] Actually the interplay between the two concepts is far tighter: see for example [Ser, Chapter IV], especially Sections A and B. is done in Example A.2.
Proposition A.6.
Let $A$ be a commutative ring and let $a_{1},\dots,a_{d}\in A$ with $I=\left(a_{1},\dots,a_{d}\right)$. Set
$s:E=A^{d}$ $\to A$
$\sum c_{i}e_{i}$ $\mapsto\sum c_{i}a_{i}$
and denote with $\bigwedge^{\bullet}\left(s\right)^{+}$ the complex
$0\to A=\bigwedge^{d}E\to\bigwedge^{d-1}E\to\dots\to E\overset{s}{\longrightarrow}A\to\frac{A}{I}\to 0.$
Then
  1. 1.
    If $a_{1},\dots,a_{d}$ is a regular sequence, then $\bigwedge^{\bullet}\left(s\right)^{+}$ is a free resolution of $A/I$.
  2. 2.
    Suppose $A$ is local, noetherian, that $a_{1},\dots,a_{d}\in\mathfrak{m}$ and that $H_{1}\left(s\right)=0$. Then $a_{1},\dots,a_{d}$ is a regular sequence and $\bigwedge^{\bullet}\left(s\right)^{+}$ is exact.
Proof.
1)
Consider $E=A^{d}$ and
$s:E$ $\to A$
$\sum c_{i}e_{i}$ $\mapsto\sum c_{i}a_{i}.$
We clearly have $H_{0}\left(s\right)=\frac{A}{I}$. We want to see, by induction on $d$, that this implies that $s$ is regular: the case $d=1$ is clear, since in that case the complex is given by
$0\to A\overset{\cdot a_{1}}{\longrightarrow}A\to 0$
with
$H_{1}\left(s\right)=\ker\left(\cdot a\right)=0$
by regularity. For greater $d$’s, looking at the exact sequence (3), we have $H_{d-1}\left(t\right)=0$ by induction assumption and hence also $H_{d}\left(s\right)=0$. Therefore we’ve seen that a regular sequence always gives rise to a regular element.
2)
Again this is clear when $d=1$, so we proceed by induction on $d$. From the exact sequence (2) we get exact
$H_{1}\left(s\right)\to H_{1}\left(\left(\bigwedge^{\bullet}\left(t\right)\right)\left[-1\right]\right)=H_{0}\left(t\right)\overset{\delta_{1}}{\longrightarrow}H_{0}\left(t\right)$
where $\delta_{1}$ is just the product by $\pm a_{d}$ on $A/\left(a_{1},\dots,a_{d-1}\right)$, therefore ur assumption that $H_{1}\left(s\right)=0$ gives that $a_{d}$ is a regular element in $A/\left(a_{1},\dots,a_{d-1}\right)$. Again by induction hypothesis, we only need to see that $H_{1}\left(t\right)=0$, but the connecting map
$H_{2}\left(\left(\bigwedge^{\bullet}\left(t\right)\right)\left[-1\right]\right)=H_{1}\left(t\right)\overset{\delta_{2}}{\longrightarrow}H_{1}\left(t\right)$
is just the product by $\pm a_{d}$ and $H_{1}\left(s\right)=0$ forces it to be surjective, but then we have $a_{d}\in\mathfrak{m}$ with $a_{d}\cdot H_{1}\left(t\right)=H_{1}\left(t\right)$. By noetherianity $H_{1}\left(t\right)$ is finitely generated and hence it vanishes by Nakayama’s lemma.
Corollary A.7.
Let $\left(A,\mathfrak{m}\right)$ be a noetherian local ring and let $a_{1},\dots,a_{d}\in\mathfrak{m}$. Then $a_{1},\dots,a_{d}$ is a regular sequence if and only if every its permutation is.
Proof.
It follows from Proposition A.6 since the condition $H_{1}\left(s\right)=0$ does not depend on the order. ∎
Lemma A.8.
Let $A$ be a commutative ring and let $a_{1},\dots,a_{d}\in A$ with $I=\left(a_{1},\dots,a_{d}\right)$. set
$s:E=A^{d}$ $\to A$
$\sum c_{i}e_{i}$ $\mapsto\sum c_{i}a_{i}.$
Then $I\cdot H_{i}\left(s\right)=0$ for every $i>0$.
Proof.
We proceed by induction on $d$, the case $d=1$ being clear. For $d>1$, in view of Remark A.5 we have exact sequences
$0\to\frac{H_{j}\left(t\right)}{a_{d}\cdot H_{j}\left(t\right)}\to H_{j}\left(s\right)\to\mathrm{Ann}_{H_{j-1}\left(t\right)}\left(a_{d}\right)\to 0$
and we conclude since both edges are annihilated by $I$. ∎
Of course everything may be globalized by patching
Definition A.9.
Given a projective variety $X$ and a locally free coherent $\mathcal{O}_{X}$-module $\mathcal{E}$ of rank $r$, for any element $s\in H^{0}\left(X,\mathcal{E}\right)$, denoting with $\mathcal{I}\subseteq\mathcal{O}_{X}$ the image of $s^{\lor}:\mathcal{E}^{\lor}\to\mathcal{O}_{X}$, we can associate a complex
$\bigwedge^{\bullet}\left(s\right):\quad 0\to\bigwedge^{r}\mathcal{E}^{\lor}\to\dots\to\mathcal{E}^{\lor}\to\mathcal{O}_{X}\to\frac{\mathcal{O}_{X}}{\mathcal{I}}\to 0.$
Proposition A.10.
Given a smooth projective variety $X$ and a locally free coherent $\mathcal{O}_{X}$-module $\mathcal{E}$ of rank $r$, for any element $s\in H^{0}\left(X,\mathcal{E}\right)$ the homology
$H_{i}\left(s\right)=\mathcal{H}_{i}\left(\bigwedge^{\bullet}\left(s\right)\right)$
has support on the zero locus $Z\left(\mathrm{Im}\left(s^{\lor}\right)\right)$.
Proof.
Everything boils down to showing that, if a sequence $a_{1},\dots,a_{d}$ in a ring $A$ gives $\left(a_{1},\dots,a_{d}\right)=A$, then the corresponding Koszul complex is exact, but this follows from Lemma A.8. ∎